OpenStudy (anonymous):

HOW DO YOU SUBTRACT THEN SIMPLIFY 41/9 - 3 5/9 CAN YOU HELP ME OUT....

OpenStudy (anonymous):

you turn the mixed fraction into a improper fraction which is 32/9. so 41-39 = 9/9 and 9/9 is 1 ^^

OpenStudy (anonymous):

do you mean: \[4 _{1/9} - 3 _{5/9}\] for \[4 _{1/9} --> (4*9+1)/9 =37/9\] for \[ 3 _{5/9} --> (3*9+5)/9 = 32/9\] so \[37/9 -32/9 = 5/9\]